# Robust Online Conformal Prediction under Uniform Label Noise

### v2

@@ 3.1 The impact of label noise on prediction sets
This bias in gradient estimation can result in two potential consequences: the conformal predictor may either fail to maintain the desired coverage or suffer from reduced efficiency (i.e., generating large prediction sets). We formalize these consequences in the following proposition:
@@
$\alpha-\frac{1}{T}\sum_{t=1}^{T}$ $\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}\leq\frac{A}{\sqrt{T}}+\frac{B}{T}$
$+\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\left|\mathcal{C}_{t}(X_{t})\right|\right]\right),$
@@
(a) CIFAR-100
(b) ImageNet
Figure 1: Performance of standard online conformal prediction under different noise rates, with ResNet18 on CIFAR-100 and ImageNet. We use noisy labels to update the threshold with a constant learning rate $\eta=0.05$.
@@
The proof is provided in Appendix C.1. In Proposition 3.1, we evaluate the long-run mis-coverage rate using $\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}$. Our analysis shows that label noise introduces a coverage gap of $\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\left|\mathcal{C}_{t}(X_{t})\right|\right]\right)$, between the actual mis-coverage rate and desired mis-coverage rate $\alpha$ (the term $\frac{A}{\sqrt{T}}+\frac{B}{T}$ will diminish and eventually approach zero, as $T$ increases). In particular, the effect of label noise on the coverage guarantee can manifest in two distinct scenarios, depending on the prediction set size:
@@
We compare the performance of online conformal prediction under different noise rates $\epsilon=0.05,0.1,0.15$, with a ResNet18 model on CIFAR-100 and ImageNet datasets. We employ LAC score (Sadinle et al., 2019) to generate prediction sets, and use noisy labels to update the threshold with a constant learning rate $\eta=0.05$. The experimental results in Figure 1 validate our theoretical analysis: label noise introduces discrepancies between the actual and target coverage rates $1-\alpha$, with higher noise rates resulting in a more pronounced coverage gap.
@@
3.2 Method
Our theoretical analysis establishes that biased gradient estimates arising from label noise can significantly impact the coverage properties of online conformal prediction. Therefore, the key challenge of the noisy setting lies in how to obtain unbiased gradient estimates without requiring ground-truth labels. In this work, we propose a novel algorithm - Noise Robust Online Conformal Prediction (dubbed NR-OCP), which updates the threshold with robust pinball loss. We begin by developing the intuition behind approximating clean pinball loss under uniform label noise.
Consider a data sample $(X,Y)$ with a noisy label $\tilde{Y}$, we denote the clean non-conformity score as $S=\mathcal{S}(X,Y)$, the noisy score as $\tilde{S}=\mathcal{S}(X,\tilde{Y})$, and the score for an arbitrary class $y\in\mathcal{Y}$ as $S_{y}=\mathcal{S}(X,y)$. Under a uniform label noise with noise rate $\epsilon\in(0,1)$, the distributions of these scores have the following relationship: $\mathbb{P}\left\{S\leq s\right\}=\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}\leq s\right\}-\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\mathbb{P}\left\{S_{y}\leq s\right\}$, for an arbitrary number $s\in\mathbb{R}$. We formally establish this equation in Lemma G.1 with a rigorous proof. This correlation motivates the following approximation:
@@ 3.2 Method
Robust pinball loss. The above decomposition suggests that the clean pinball loss can be approximated by replacing its indicator function with the above expression. Inspired by this, we propose the robust pinball loss as:
$\tilde{l}_{1-\alpha}(\tau,\tilde{S},\{S_{y}\}_{y=1}^{K})=l_{1}(\tau,\tilde{S})-l_{2}(\tau,\{S_{y}\}_{y=1}^{K}),$ (3)
where
$l_{1}(\tau,\tilde{S})=\frac{1}{1-\epsilon}l_{1-\alpha}(\tau,\tilde{S}),$
$l_{2}(\tau,\{S_{y}\}_{y=1}^{K})=\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}l_{1-\alpha}(\tau,S_{y}).$
@@
The proof can be found in Appendix C.2. In Proposition 3.2, the first property ensures that our robust pinball loss matches the expected value of the true pinball loss, while the second guarantees that the gradients of both losses have the same expectation. These properties establish that updating the threshold with robust pinball loss is equivalent to updating with clean pinball loss in expectation.
With the robust pinball loss, the proposed NR-OCP updates the threshold as follows:
@@
$=\hat{\tau}_{t}-\eta\cdot\nabla_{\hat{\tau}_{t}}l_{1}(\hat{\tau}_{t},\tilde{S}_{t})+\eta\cdot\nabla_{\hat{\tau}_{t}}l_{2}(\hat{\tau}_{t},\{S_{t,y}\}_{y=1}^{K}),$
where
$\nabla_{\hat{\tau}_{t}}l_{1}(\hat{\tau}_{t},\tilde{S}_{t})=\frac{1}{1-\epsilon}\left[\mathds{1}\left\{\tilde{S}_{t}\leq\hat{\tau}_{t}\right\}-(1-\alpha)\right],$
$\nabla_{\hat{\tau}_{t}}l_{2}(\hat{\tau}_{t},S_{t,y})=\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\left[\mathds{1}\left\{S_{t,y}\leq\hat{\tau}_{t}\right\}-(1-\alpha)\right].$
In summary, Proposition 3.2 establishes the validity of our gradient estimates in expectation, laying the foundation for our subsequent convergence analysis. In the following sections, we present a detailed theoretical analysis that demonstrates how these results translate into finite-sample coverage guarantees, in the presence of label noise.
3.3 Convergence of coverage rate
We now analyze the convergence of the coverage rate of NR-OCP under uniform label noise with a constant learning rate schedule. For notation shorthand, we denote:
$\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}\right]=\mathbb{E}_{S_{t}}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},S_{t})\right],$
$\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}\right]=\mathbb{E}_{\tilde{S}_{t},S_{t,y}}\left[\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}(\hat{\tau}_{t},\tilde{S}_{t},\{S_{t,y}\}_{y=1}^{K})\right].$
@@ 3.3 Convergence of coverage rate
The proof of Proposition 3.3 (in Appendix C.3) relies on the following decomposition:
$\quad\mathrm{ExErr(T)}=\left|\sum_{t=1}^{T}\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}\right]\right|=\left|\sum_{t=1}^{T}\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}\right]\right|$
$\leq\underbrace{\left|\sum_{t=1}^{T}\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}\right]-\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}\right|}_{(a)}+\underbrace{\left|\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}\right|}_{(b)}.$
@@
Building on Proposition 3.3, we now provide an upper bound for empirical coverage error for NR-OCP:
@@
$\mathrm{EmErr(T)}\leq\,\frac{2-\epsilon}{1-\epsilon}\sqrt{2\log\left(\frac{4}{\delta}\right)}\cdot\frac{1}{\sqrt{T}}+$
$\left(\frac{1}{\eta}-\alpha+\frac{\epsilon}{1-\epsilon}\right)\cdot\frac{1}{T}.$
@@
The proof of Proposition 3.4 (detailed in Appendix C.4) employs a similar decomposition:
$\quad\mathrm{EmErr(T)}=\left|\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}l_{1-\alpha}\right|$
$\leq\underbrace{\left|\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}l_{1-\alpha}-\sum_{t=1}^{T}\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}\right]\right|}_{(a)}+\underbrace{\left|\sum_{t=1}^{T}\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}\right]\right|}_{(b)}.$
The analysis shows that both terms achieve $\mathcal{O}(T^{-1/2})$ convergence: part (a) through the Azuma–Hoeffding inequality, and part (b) via Proposition 3.3. Combining the two parts yields the desired upper bound.
@@
It is worth noting that our method achieves a $\mathcal{O}(T^{-1/2})$ convergence rate for empirical coverage error even in the absence of noise (i.e., $\epsilon=0$), which is slightly slower than the $\mathcal{O}(T^{-1})$ rate achieved by standard online conformal prediction theory (Gibbs & Candes, 2021; Angelopoulos et al., 2024). This is because our analysis relies on martingale-based concentration to handle label noise, leading to the $\mathcal{O}(T^{-1/2})$ rate.
4 Results for dynamic learning rate
4.1 Theoretical results
Recent work (Angelopoulos et al., 2024) highlights a limitation of constant learning rates: while coverage holds on average over time, the instantaneous coverage rate $\mathrm{Cov}(\hat{\tau}_{t})=\mathbb{P}\left\{S\leq\hat{\tau}_{t}\right\}$ would exhibit substantial temporal variability (see Proposition 1 in Angelopoulos et al. (2024)). Thus, they propose using dynamic learning rates where $\eta_{t}$ can change over time for updating the threshold.
We analyze this setting in the presence of label noise, specifically examining the case where noisy labels update the conformal threshold with a time-varying learning rate:
$\hat{\tau}_{t+1}=\hat{\tau}_{t}-\eta_{t}\cdot\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},\tilde{S}_{t}).$ (5)
For notational simplicity, we denote $\eta_{0}^{-1}=0$. Then, we have the following proposition analogous to Proposition 3.1:
@@ 4.1 Theoretical results
$\alpha-\frac{1}{T}\sum_{t=1}^{T}$ $\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}\leq\frac{A}{\sqrt{T}}+\frac{B}{T}$
$+\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\left|\mathcal{C}_{t}(X_{t})\right|\right]\right)$
where
@@
$A=\frac{2-\epsilon}{1-\epsilon}\sqrt{2\log\left(\frac{4}{\delta}\right)};$
$B=\frac{1+\underset{1\leq t\leq T-1}{\max}\eta_{t}}{1-\epsilon}\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|.$

### v3

@@ 3 The impact of label noise on online conformal prediction
This bias in gradient estimation can result in two potential consequences: conformal predictors may either fail to maintain the desired coverage or suffer from reduced efficiency (i.e., generating large prediction sets). We formalize these consequences in the following proposition:
@@
Consider online conformal prediction under uniform label noise with noise rate $\epsilon\in(0,1)$. Given Assumptions 2.1 and 2.2, when updating the threshold according to Eq. (3), then for any $\delta\in(0,1)$ and $T\in\mathbb{N}^{+}$, the following bound holds with probability at least $1-\delta$:
$\alpha-\frac{1}{T}\sum_{t=1}^{T}$ $\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}\leq\frac{A}{\sqrt{T}}+\frac{B}{T}+\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\left|\mathcal{C}_{t}(X_{t})\right|\right]\right),$
@@
The proof is provided in Appendix E.1. In Proposition 3.1, we evaluate the long-run mis-coverage rate using $\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}$. Our analysis shows that label noise introduces a coverage gap of
$\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\left|\mathcal{C}_{t}(X_{t})\right|\right]\right),$
between the actual mis-coverage rate and desired mis-coverage rate $\alpha$ (the term $\frac{A}{\sqrt{T}}+\frac{B}{T}$ will diminish and eventually approach zero, as $T$ increases). In particular, the effect of label noise on the coverage guarantee can manifest in two distinct scenarios, depending on the prediction set size:
@@
To verify our theoretical result, we compare the performance of ACI under different noise rates. In particular, we compute the long-run coverage, defined as $\mathrm{Cov}(T)=\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\in\mathcal{C}_{t}(X_{t})\right\}$, under noise rate $\epsilon\in\{0.05,0.1,0.15\}$, with a ResNet18 model on CIFAR-100 and ImageNet datasets. We employ LAC score (sadinle2019least, ) to generate prediction sets with error rate $\alpha=0.1$, and use noisy labels to update the threshold with a constant learning rate $\eta=0.05$. The experimental results in Figure 1 validate our theoretical analysis: label noise introduces discrepancies between the actual and target coverage rates $1-\alpha$, with higher noise rates resulting in a more pronounced coverage gap.
@@
(a) CIFAR-100
(b) ImageNet
Figure 1: Performance of ACI (gibbs2021adaptive, ) with a constant learning rate $\eta=0.05$ under different noise rates, using ResNet18 on CIFAR-100 and ImageNet datasets. The results validate that label noise introduces a coverage gap, with higher noise rates resulting in a more pronounced gap.
4 Method
4.1 Robust pinball loss
Our theoretical analysis establishes that biased gradient estimates arising from label noise can significantly impact the coverage properties of online conformal prediction. Therefore, the key challenge of the noisy setting lies in how to obtain unbiased gradient estimates without requiring ground-truth labels. In this work, we propose a novel loss function - robust pinball loss, which provides an unbiased estimate of the clean pinball loss value without requiring access to ground-truth labels. We begin by developing the intuition behind how to approximate the clean pinball loss under uniform label noise.
[…]
@@ 4.1 Robust pinball loss
The above decomposition suggests that the clean pinball loss can be approximated by replacing its indicator function with the above expression. Inspired by this, we propose the robust pinball loss as:
$\tilde{l}_{1-\alpha}(\tau,\tilde{S},\{S_{y}\}_{y=1}^{K})=\frac{1}{1-\epsilon}l_{1-\alpha}(\tau,\tilde{S})-\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}l_{1-\alpha}(\tau,S_{y}).$ (4)
@@
The proof can be found in Appendix E.2. In Proposition 4.1, the first property ensures that our robust pinball loss matches the expected value of the true pinball loss, while the second guarantees that the gradients of both losses have the same expectation. These properties establish that updating the threshold with robust pinball loss is equivalent to updating with clean pinball loss in expectation.
In summary, Proposition 4.1 establishes the validity of our robust pinball loss in expectation. Notably, our robust pinball loss offers a general solution to the uniform label noise, and is complementary to existing online conformal prediction methods. In the following sections, we apply the robust pinball loss to ACI with constant (gibbs2021adaptive, ) and dynamic learning rates (DBLP:conf/icml/AngelopoulosBB24, ). We will show that by updating the threshold with the proposed loss, ACI eliminates the coverage gap caused by the label noise.
4.2 Convergence with constant learning rate
We now analyze the convergence of the coverage rate of ACI under uniform label noise with a constant learning rate schedule (gibbs2021adaptive, ). In particular, we update the threshold of ACI with respect to the robust pinball loss:
@@ 4.2 Convergence with constant learning rate
$=\hat{\tau}_{t}-\eta\cdot\frac{1}{1-\epsilon}\left[\mathds{1}\left\{\tilde{S}_{t}\leq\hat{\tau}_{t}\right\}-(1-\alpha)\right]+\eta\cdot\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\left[\mathds{1}\left\{S_{t,y}\leq\hat{\tau}_{t}\right\}-(1-\alpha)\right].$
For notation shorthand, we denote:
$\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}\right]=\mathbb{E}_{S_{t}}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},S_{t})\right],$ $\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}\right]=\mathbb{E}_{\tilde{S}_{t},S_{t,y}}\left[\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}(\hat{\tau}_{t},\tilde{S}_{t},\{S_{t,y}\}_{y=1}^{K})\right].$
@@
Under the same assumptions in Proposition 3.1, when updating the threshold according to Eq. (5), then for any $\delta\in(0,1)$ and $T\in\mathbb{N}^{+}$, the following bound holds with probability at least $1-\delta$:
@@
The proof (in Appendix E.3) relies on the following decomposition:
$\quad\mathrm{ExErr(T)}=\left|\sum_{t=1}^{T}\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}\right]\right|\leq\underbrace{\left|\sum_{t=1}^{T}\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}\right]-\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}\right|}_{(a)}+\underbrace{\left|\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}\right|}_{(b)}.$
@@
Building on Proposition 4.2, we now provide an upper bound for the empirical coverage error:
@@
Under the same assumptions in Proposition 3.1, when updating the threshold according to Eq. (5), then for any $\delta\in(0,1)$ and $T\in\mathbb{N}^{+}$, the following bound holds with probability at least $1-\delta$:
@@
The proof (detailed in Appendix E.4) employs a similar decomposition:
$\quad\mathrm{EmErr(T)}=\left|\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}l_{1-\alpha}\right|\leq\underbrace{\left|\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}l_{1-\alpha}-\sum_{t=1}^{T}\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}\right]\right|}_{(a)}+\underbrace{\left|\sum_{t=1}^{T}\mathbb{E}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}\right]\right|}_{(b)}.$
The analysis shows that both terms achieve $\mathcal{O}(T^{-1/2})$ convergence: part (a) through the Azuma–Hoeffding inequality, and part (b) via Proposition 4.2.
@@
It is worth noting that our method achieves a $\mathcal{O}(T^{-1/2})$ convergence rate for empirical coverage error even in the absence of noise (i.e., $\epsilon=0$), which is slightly slower than the $\mathcal{O}(T^{-1})$ rate achieved by standard online conformal prediction theory (gibbs2021adaptive, ; DBLP:conf/icml/AngelopoulosBB24, ). This is because our analysis relies on martingale-based concentration to handle label noise, leading to the $\mathcal{O}(T^{-1/2})$ rate.
4.3 Convergence with dynamic learning rate
Recent work (DBLP:conf/icml/AngelopoulosBB24, ) highlights a limitation of constant learning rates: while coverage holds on average over time, the instantaneous coverage rate $\mathrm{Cov}(\hat{\tau}_{t})=\mathbb{P}\left\{S\leq\hat{\tau}_{t}\right\}$ would exhibit substantial temporal variability (see Proposition 1 in (DBLP:conf/icml/AngelopoulosBB24, )). Thus, they extend ACI to dynamic learning rate schedule where $\eta_{t}$ can change over time for updating the threshold. In this section, we apply the robust pinball loss to the dynamic learning rate schedule. Specifically, we update the threshold by:
$=\hat{\tau}_{t}-\eta_{t}\cdot\frac{1}{1-\epsilon}\left[\mathds{1}\left\{\tilde{S}_{t}\leq\hat{\tau}_{t}\right\}-(1-\alpha)\right]+\eta_{t}\cdot\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\left[\mathds{1}\left\{S_{t,y}\leq\hat{\tau}_{t}\right\}-(1-\alpha)\right].$
For the convergence of coverage rate, our theoretical results show that, under dynamic learning rate, NR-OCP achieves convergence rates of $\mathrm{EmCovErr(T)}=\mathcal{O}(T^{-1/2})$ and $\mathrm{ExCovErr(T)}=\mathcal{O}(T^{-1/2})$. The proofs are presented in Appendix E.5 and E.6.
@@ 4.3 Convergence with dynamic learning rate
Under the same assumptions in Proposition 3.1, when updating the threshold according to Eq. (6), then for any $\delta\in(0,1)$ and $T\in\mathbb{N}^{+}$, the following bound holds with probability at least $1-\delta$:
Proposition 4.6.
Under the same assumptions in Proposition 3.1, when updating the threshold according to Eq. (6), then for any $\delta\in(0,1)$ and $T\in\mathbb{N}^{+}$, the following bound holds with probability at least $1-\delta$:
@@
$\mathrm{EmErr(T)}\leq\frac{2-\epsilon}{1-\epsilon}\sqrt{2\log\left(\frac{4}{\delta}\right)}\cdot\frac{1}{\sqrt{T}}+\left[\left(1+\max_{1\leq t\leq T-1}\eta_{t}\cdot\frac{1+\epsilon}{1-\epsilon}\right)\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|\right]\cdot\frac{1}{T}.$
In Proposition 4.5 and 4.6, we establish that for any sequence of learning rates that satisfies the condition $\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|/T\to 0$ as $T\to+\infty$, both expected and empirical coverage errors asymptotically vanish with a convergence rate of $\mathcal{O}(T^{-1/2})$. Therefore, by applying our method, the long-term coverage rate would approach the desired target level of $1-\alpha$.
Online learning theory (see Theorem 2.13 of (orabona2019modern, )) established that ACI achieves a regret bound:
$\mathrm{Reg}(T)=\sum_{t=1}^{T}l_{1-\alpha}(\hat{\tau}_{t},S_{t})-\min_{\tau}\sum_{t=1}^{T}l_{1-\alpha}(\tau,S_{t})=\mathcal{O}(T^{-\frac{1}{2}}),$
with optimally chosen $\eta_{t}$, which serves as a helpful measure alongside coverage (bhatnagar2023improved, ). We provide a regret analysis for our method in Appendix B.1. In addition, we formally establish the convergence of $\hat{\tau}_{t}$ towards the global minima in Appendix B.2 by extending the standard convergence analysis of stochastic gradient descent (see Theorem 5.3 in (garrigos2023handbook, )).
Figure 2: Long-run coverage and local coverage performance of various methods under uniform noisy labels with noise rate $\epsilon=0.05$. We apply robust pinball loss to ACI with (a) constant (gibbs2021adaptive, ) and (b) dynamic learning rate (DBLP:conf/icml/AngelopoulosBB24, ), and (c) SAOCP (bhatnagar2023improved, ). We employ LAC to generate prediction sets with $\alpha=0.1$, using ResNet18 on CIFAR100. “Baseline” and “Clean” denote the online conformal prediction with standard pinball loss, using noisy and clean labels.
5 Experiments
5.1 Experimental setups
Datasets and setup. We evaluate the performance of NR-OCP under uniform label noise. The experiments include both constant $\eta=0.05$ and dynamic learning rates $\eta_{t}=1/t^{1/2+\varepsilon}$ with $\varepsilon=0.1$, following prior work (DBLP:conf/icml/AngelopoulosBB24, )). We use CIFAR-100 (krizhevsky2009learning, ) and ImageNet (deng2009imagenet, ) datasets with synthetic label noise. On ImageNet, we use four pre-trained classifiers from TorchVision (paszke2019pytorch, ) - ResNet18, ResNet50 (he2016deep, ), DenseNet121 (huang2017densely, ) and VGG16 (simonyan2014very, ). On CIFAR-100, we train these models for 200 epochs using SGD with a momentum of 0.9, a weight decay of 0.0005, and a batch size of 128. We set the initial learning rate as 0.1, and reduce it by a factor of 5 at 60, 120 and 160 epochs.
Conformal prediction algorithms. We integrate the robust pinball loss into ACI with constant (gibbs2021adaptive, ) and dynamic learning rates (DBLP:conf/icml/AngelopoulosBB24, ), and strongly adaptive online conformal prediction (bhatnagar2023improved, ). We apply LAC (sadinle2019least, ), APS (romano2020classification, ), RAPS (angelopoulos2020uncertainty, ) and SAPS (huang2024uncertainty, ) to generate prediction sets. A detailed description of these non-conformity scores is provided in Appendix B.4. In addition, we use error rates $\alpha\in\{0.1,0.05\}$ in experiments.
Metrics. We employ four evaluation metrics: long-run coverage (Cov), local coverage (LocalCov), coverage gap (CovGap) and prediction set size (Size). We use long-run coverage and local coverage to present the dynamics of coverage during the optimization. Formally, for any $T\in\mathbb{N}^{+}$,
$\mathrm{Cov}(T)=\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\in\mathcal{C}_{t}(X_{t})\right\},\quad\mathrm{LocalCov}(T)=\frac{1}{L}\sum_{t=T}^{T+L}\mathds{1}\left\{Y_{t}\in\mathcal{C}_{t}(X_{t})\right\},$
In particular, long-run coverage measures the coverage rate over the first $T$ time steps, while local coverage is over an interval with length $L$. In the experiments, we employ a length of $L=200$. Besides, the coverage gap and the prediction set size are computed over the full test set, indicating the final performance of the online conformal prediction method. Formally, given a test dataset $\mathcal{I}_{test}$,
$\mathrm{CovGap}=\left|\frac{1}{|\mathcal{I}_{test}|}\sum_{t\in\mathcal{I}_{test}}\mathds{1}\left\{Y_{t}\in\mathcal{C}_{t}(X_{t})\right\}-(1-\alpha)\right|,\quad\mathrm{Size}=\frac{1}{|\mathcal{I}_{test}|}\sum_{t\in\mathcal{I}_{test}}|\mathcal{C}_{t}(X_{t})|.$
Small prediction sets are preferred as they can provide precise predictions, thereby enabling accurate human decision-making in real-world scenarios (DBLP:conf/icml/CresswellSKV24, ).
5.2 Main results
Robust pinball loss enhances the noise robustness of existing online conformal prediction methods. In figure 2, we apply robust pinball loss to ACI with constant (gibbs2021adaptive, ) and dynamic learning rate (DBLP:conf/icml/AngelopoulosBB24, ), and SAOCP (bhatnagar2023improved, ). We evaluate these methods with long-run coverage and local coverage. We use the LAC score and a ResNet18 model on CIFAR100, with error rate $\alpha=0.1$ and noise rate $\epsilon=0.05$. The results demonstrate that the proposed robust pinball loss enables these methods to achieve a long-run coverage and local coverage close to the target coverage rate of $1-\alpha$. In summary, the empirical results highlight the versatility of our robust pinball loss, demonstrating its potential to enhance the noise robustness across diverse algorithms.
Robust pinball loss is effective across various settings. In Table 1, we incorporate robust pinball loss into ACI with constant (gibbs2021adaptive, ) and dynamic learning rate (DBLP:conf/icml/AngelopoulosBB24, ), under different noise rate $\epsilon$, error rate $\alpha$ and non-conformity scores. We evaluate the prediction sets with coverage gap and prediction set size. Due to space constraints, we report the average performance across four non-conformity score functions. The performance on each score function is provided in Appendix B.5. The results show that robust pinball loss allows ACI to eliminate the coverage gap, achieving precise coverage guarantees while significantly improving the long-run efficiency of prediction sets. For example, on ImageNet with error rate $\alpha=0.1$, noise rate $\epsilon=0.15$ and dynamic learning rate, ACI employing the robust pinball loss achieves a negligible coverage gap of 0.183% and a prediction set size of 13.10. We report additional results on various models in Appendix B.6. Overall, the robust pinball loss is effective across different settings, including various label noise, error rate, non-conformity score functions, and model architectures.
In real-world scenarios where the noise rate could be unknown, it can be estimated from historical data liu2015classification ; yu2018efficient ; wei2020combating . In Appendix C, we leverage an existing approach to estimate the noise rate. Then, we employ this estimated noise rate in our method. We show that in this circumstance, the robust pinball loss can improve the noise robustness of online conformal prediction algorithms.